OpenStudy (anonymous):

How many grams of magnesium metal will react completely with 6.3 liters of 4.5 M HCl? Show all of the work needed to solve this problem. Mg (s) + 2 HCl (aq)yields MgCl2 (aq) + H2 (g)

OpenStudy (aaronq):

you need to find the moles present of HCl. Use: \(molarity=\dfrac{moles}{L_{solution}}\) Next relate the moles and stoichiometric coefficients to find the moles of Mg \(\sf \dfrac{moles~of~Mg}{Mg's~coefficient}=\dfrac{moles~of~HCl}{HCl's~coefficient}\) next convert moles of Mg to mass \(moles=\dfrac{moles}{Molar~mass}\)